# SynPO: Synergizing Descriptiveness and Preference Optimization for Video Detailed Captioning

### v1

@@ 1 Introduction
Fine-grained video captioning aims to generate detailed and coherent textual descriptions that precisely capture video contents. This task necessitates the recognition of salient actions and objects, while also modeling fine-grained visual features and temporal dynamics. Recent studies [cheng2024videollama2advancingspatialtemporal, jin2024chat, cai2024matryoshka, li2024llama] have primarily employed Vision-Language Models (VLMs) for video captioning. These methods [li2023blip, li2023videochat, zhang2023video] typically utilize pre-trained vision encoders and Large Language Models (LLMs), with a connector module linking them. By training on video-text pairs, these models aim to align visual and textual representations, thereby enhancing their ability to understand and describe video content [wang2023chatvideo] effectively.
@@
Direct Preference Optimization (DPO) [rafailov2024direct] is a fine-tuning method that aligns models with stipulated preferences using high-quality preference pairs. Recent work has successfully adapted DPO and its variants to video understanding tasks, significantly improving model performance [li2025templetemporalpreferencelearningvideo, li2025temporalpreferenceoptimizationlongform]. Thus, integrating DPO into fine-grained video captioning to enhance the model’s ability to capture temporal dynamics and detailed descriptions present a promising direction. However, two critical challenges currently degrade its performance in fine-grained video captioning: (1) the scarcity of high-quality video-text alignment pairs, which are essential for preference learning; (2) DPO typically suffers from the simultaneous decrease in both positive and negative reward values [pal2024smaugfixingfailuremodes, chen2024improvedpreferenceoptimizationpipeline], leading to a potential objective optimization deviation from focusing on generation quality to merely discriminating between preferences [feng2024analyzingunderstandinglimitationsdpo], as shown in Figure 1 (middle).
@@
Recently, the proposed VDC benchmark [chai2024auroracap] is well-curated, but limited in scale and partially reliant on manual annotations. Other video captioning datasets, such as MSRVTT [xu2016msr], VATEX [wang2020vatexlargescalehighqualitymultilingual], MSVC [cheng2024videollama2advancingspatialtemporal], etc., typically provide overly brief captions, falling short in fine-grained video captioning. Besides, these datasets lack preference pairs, and thus cannot be prepared for DPO. To construct preference pairs, many existing methods [an2025agfsyncleveragingaigeneratedfeedback, Li_2024] rely on a stronger VLM to score multiple outputs from the same prompt. While straightforward, this approach is impractical: small teams face prohibitive API costs, while developers of powerful models often lack access to a stronger scoring model. Some studies attempt to circumvent this limitation by generating negative preferences, such as through atypical item substitution [xie2024vdpomitigatinghallucinationlarge] or temporal perturbation [li2025templetemporalpreferencelearningvideo]. However, these methods primarily focus on negative samples and fail to produce higher-quality positive preferences.
@@
We present an automated pipeline for constructing high-quality preference pairs for fine-grained video captioning. Given the same input, we generate multiple alternative outputs using a VLM. These candidates are then scored leveraging intrinsic properties of the VLM itself, such as its self-consistency [10298561] and the enhanced ability to capture details in short videos, with limited assistance from an LLM. The top and bottom scores are selected as positive and negative preferences, respectively, forming our constructed preference dataset. Compared to existing approaches, our method achieves an optimal balance between cost efficiency and high-quality preference pair construction.
@@
Our experiments across multiple models and datasets demonstrate that our data construction pipeline can generate high-quality preference datasets with considerable generality. In addition to video captioning benchmarks (e.g., VDC [chai2024auroracap], VDD [li2023videochat], VATEX [wang2020vatexlargescalehighqualitymultilingual], MSR-VTT [xu2016msr]), we conduct comparisons of SynPO with various DPO variants [Ethayarajh2024KTOMA, xu2024contrastive, pal2024smaugfixingfailuremodes, Azar2023AGT, meng2024simposimplepreferenceoptimization] across multiple NLP tasks, which include preference evaluation tasks (e.g., MT-Bench [zheng2023judging], AlpacaEval2 [li2023alpacaeval]) and downstream applications (e.g., tasks from the Huggingface Open LLM Leaderboard [open-llm-leaderboard-v1, open-llm-leaderboard-v2]). As shown in Figure 1 (left), extensive results indicate that SynPO significantly outperforms DPO and its variants.
@@
Figure 1: Left: SynPO significantly outperforms other methods in different models on VDC benchmark [chai2024auroracap]. Middle: Language capability degradation occurs during the latter training stages in DPO. Training collapses and is biased towards maximazing positive-negative reward gap. Right: SynPO mitigates degradtion successfully and resolves the issue of optimization objectives shifting from language capability to ranking differentiation. Its performance significantly outperforms that of DPO.
@@ 2 Related Work
Video Captioning. Early video captioning methods employed template-based approaches [guadarrama2013youtube2text] or RNN-based encoder-decoder frameworks [venugopalan2014translating], but were limited in modeling long-range dependencies. Transformers [vaswani2017attention] [dao2022flashattentionfastmemoryefficientexact] [jin2024mohmultiheadattentionmixtureofhead] [dai2019transformerxlattentivelanguagemodels] [parmar2018imagetransformer] and vision-language pre-training significantly advanced the field, with models like CLIP4Caption [tang2021clip4caption] and SwinBERT [lin2022swinbert] improving video-text alignment. Recent VLMs [cheng2024videollama2advancingspatialtemporal, li2024llama] adapt multimodal architectures such as BLIP-2 [li2023blip] and LLaVA [liu2023visual] to video, though many struggle with temporal dynamics. Newer approaches like VideoLLaMA [zhang2023video] and ChatVideo [wang2023chatvideo] better model sequential structure for enhanced comprehension.
Reinforcement Learning and Preference Learning. Reinforcement Learning from Human Feedback (RLHF) [christiano2017deep] aligns LLMs with human preferences [kaufmann2024surveyreinforcementlearninghuman, Ouyang2022TrainingLM, stiennon2020learning] through supervised fine-tuning [zhou2024lima], reward modeling [gao2023scaling], and policy optimization, improving instruction-following [ouyang2022training] and safety [bai2022training]. To simplify RLHF’s complexity, offline methods like DPO [rafailov2024direct] bypass explicit reward modeling, inspiring variants such as IPO [azar2024general], ORPO [hong2024orpo] and others [yuan2024rrhf, xucontrastive, xiao2024leverage, wu2024alphadpoadaptiverewardmargin].
@@ 3.1 Enhanced Model Inference
Contrastive decoding was initially proposed by [leng2023mitigatingobjecthallucinationslarge] to reduce object hallucinations and subsequently improved by [zhang2025eventhallusiondiagnosingeventhallucinations], who introduced a more efficient variant that contrasts logits from sparse frame samples with those from full sequences. In this work, the improved contrastive decoding method is adopted to suppress overconfidence in noisy or irrelevant features, thereby improving factual consistency and detail accuracy in generated captions.
The self-retrospective strategy, proposed by [ahn2025isrdpoaligninglargemultimodal], operates outside the decoding process and enhances comprehension through iterative refinement. Feeding the model’s own outputs back into its input enables a form of retrospective reasoning [madaan2023selfrefineiterativerefinementselffeedback], allowing predictions to be refined in light of prior generations. We adapt this method to video captioning by using the initial caption as contextual input for subsequent refinement steps, enabling richer and more coherent descriptions.
@@ 3.2 Dataset Construction Pipeline
We propose an automated pipeline for constructing high-quality preference pairs specifically designed for fine-grained video captioning. The overall framework is designed to address the limitations of existing datasets, such as limited scale, insufficient detail in captions, and the lack of human-like preference annotations [chai2024auroracap, wang2020vatexlargescalehighqualitymultilingual, cheng2024videollama2advancingspatialtemporal]. Our method leverages both the intrinsic properties of VLMs and the reasoning capabilities of LLMs to generate diverse and reliable preference pairs without dependence on costly multimodal scorers.
The core strategy involves generating multiple candidate captions per video using a single VLM under the same prompt. These candidates are then scored using a novel three criterion evaluation framework that combines factuality, modality correctness, and self-consistency [10298561]. Based on the aggregated scores across all three criteria, the captions with the highest and lowest total scores are selected as positive and negative preferences, respectively, forming our final dataset. In the following, we describe each scoring criterion in detail (full prompts provided in Appendix D).
Criterion 1: Factuality through Temporal Decomposition. Due to input length limits in most VLMs, processing long videos directly often causes detail loss and hallucinations [li2025templetemporalpreferencelearningvideo]. To mitigate this, we divide each video into short clips, process them independently with the VLM to generate clip-level captions, and concatenate these into a reference set. An LLM then assesses the consistency between the full-video caption and the reference set, focusing particularly on factual alignment. This approach enhances detail preservation and mitigates hallucinations. Scores range from 0 to 5.
@@
Criterion 3: Self-consistency through Multi-sample Analysis. Inspired by self-consistency methods in NLP [rueda2025understandingllmscientificreasoning, liu2025enhancingmathematicalreasoninglarge, sanzguerrero2025correctiveincontextlearningevaluating], we apply it to video captioning by assessing the stability of key entities, actions, and temporal dynamics across multiple generations. Specifically, the VLM generates $n$ diverse captions via high-temperature sampling. An LLM analyzes their similarity, rewarding consistent patterns and penalizing outliers through a majority voting mechanism. Given its narrower discriminative capacity compared to the other two criteria, this metric uses a 0 to 3 scoring range.
@@ 4.1 Preliminary
In contrast, DPO [rafailov2024direct] introduces a simplified framework for preference optimization that bypasses explicit reward modeling. It directly formulates the preference optimization problem as a classification task over preference pairs, eliminating the need for a separate reward model. Compared with RLHF, DPO simplifies the training pipeline while achieving competitive performance in multiple tasks such as dialogue generation. Its objective function is defined as:
@@ 4.2.1 Existing Limitations
Despite practical success of DPO, several studies have identified limitations within DPO. Specifically, [pal2024smaugfixingfailuremodes] point out that the DPO loss depends solely on the difference between the log-probability ratios of positive and negative preferences. Such a manner suggests that the final loss can decrease even when both are reduced, as long as the negative response decreases more rapidly. [meng2024simposimplepreferenceoptimization] observe preference optimization algorithms tend to decrease downstream task performance. Moreover, [xiao2025simperminimalistapproachpreference] show that DPO frequently leads to imbalanced updates between positive and negative preferences. As illustrated in Figure 3, this imbalance manifests as a concurrent decline in both positive and negative rewards during training, particularly under higher learning rates.
@@ 4.2.2 Theoretical Insights
We argue that this substitution in the DPO derivation induces a fundamental deviation in the model’s optimization objective. In DPO, minimizing the loss is equated with improving the model’s ability to rank positive preferences above negative ones, rather than generating higher-quality outputs. Consequently, the model may behave more like a ranking model than a generative one. This deviation from the original goal of RLHF, namely, generating high-reward coherent text, can lead to suboptimal outcomes in terms of language capability. Furthermore, the logarithmic term derived from the KL-divergence constraint in Eq. (2) fails to serve its intended role in DPO. Due to the derivative properties of the logarithmic function, decreasing reward values require smaller gradient steps than increasing them [xiao2025simperminimalistapproachpreference]. As a result, the optimizer is incentivized to reduce both positive and negative rewards simultaneously to minimize the overall loss.
@@ 4.2.3 Empirical Observations
To further investigate this behavior, we analyze the gradient dynamics of the DPO loss with respect to the policy parameters. According to the original DPO paper [rafailov2024direct], the gradient is derived as:
@@
However, empirical success has been reported in the original DPO paper and its variants (e.g., DPOP [pal2024smaugfixingfailuremodes], IPO [Azar2023AGT]).These results appear inconsistent with our theoretical findings on the limitations of DPO-style optimization. To understand this discrepancy, we analyze their experimental setups and identified two key factors contributing to the observed performance improvements:
(1) Low Learning Rates Mitigate Instability. Most DPO-style methods use significantly lower learning rates compared to standard Supervised Fine-Tuning (SFT). As shown in the original DPO paper [rafailov2024direct], a learning rate of 1e-6 is used—much lower than the typical SFT setting of 2e-5 (other variant configurations provided in Tabel 8). Our gradient-based analysis reveals that under such settings, the magnitude of parameter updates is less than one-tenth of that in standard SFT. This implicitly constrains the model’s deviation from its initial state, thereby alleviating the negative effects arising from the deviation of the optimization objective.
@@ 5.1 Constructing Long Video Caption Preference Pairs
Enhanced Inference Evaluation. The impact of contrastive decoding and the self-retrospective strategy on the quality of generated captions is evaluated. We compare four settings: baseline (no enhancement), contrastive decoding only, self-retrospective only, and the combination of both. Each setting is applied to generate captions on the dataset of VDD [li2023videochat], and the outputs are evaluated by an LLM along six dimensions: accuracy, richness, completeness, fluency, dynamics and coherence (full prompts provided in Appendix D). As shown in Table 1, combining both strategies achieves the highest scores across all metrics, outperforming either method alone. Specifically: (1) Contrastive decoding significantly improves accuracy (+6.7%) and completeness (+2.8%), indicating its effectiveness in reducing hallucinations and ensuring factual consistency; (2) The self-retrospective strategy excels in enhancing richness (+5.5%) and dynamics (+7.6%), demonstrating its ability to inject more detailed content; (3) The combined approach retains the strengths of both methods, achieving balanced improvements in all aspects.
@@
To assess the contribution of each of our three proposed scoring criteria to the quality of preference pairs, we conduct ablation studies using AuroraCap [chai2024auroracap] as the base model, generating 10 samples per input with the self-retrospective strategy. Results for different criterion combinations are reported in Table 2. The results show that all three criteria meaningfully contribute
@@
Downstream Preference Learning. To validate the effectiveness of the proposed pipeline, AuroraCap is fine-tuned using several methods, including SFT, DPO [rafailov2024direct], and our SynPO. The preference pairs used for fine-tuning were generated from a subset of Sharegpt4video [chen2024sharegpt4video] through our automated pipeline, with additonal details provided in Appendix C. Experiments are further conducted across diverse datasets and model configurations. The resulting models are evaluated on standard video captioning benchmarks (MSR-VTT [xu2016msr], VATEX [wang2020vatexlargescalehighqualitymultilingual]) using CIDEr (C) [vedantam2015cider] and METEOR (M) [banerjee2005meteor] metrics. In addition, we assess performance on VDC [chai2024auroracap] and VDD [li2023videochat] benchmarks, which employ LLM-based evaluation to better measure linguistic richness and factual consistency through their longer, richer reference captions(see Appendix C for details). As shown in Table 3, all fine-tuned models significantly surpassed their baselines, confirming that our data construction pipeline yields high-quality data.
Table 3: Experimental evaluation on various models (AuroraCap [chai2024auroracap], LLaVA1.6-7B-video [liu2024llavanext], InterVL2-8B [wang2025enhancingreasoningabilitymultimodal]), datasets (Sharegpt4video [chen2024sharegpt4video], Charades [sigurdsson2016hollywoodhomescrowdsourcingdata], Pandas-70M [chen2024panda70mcaptioning70mvideos]), and fine-tuning approaches (SFT, DPO, SynPO and other variants).
@@
AuroraCap (DPO [rafailov2024direct]) 1.39 1.89 1.70 1.73 1.94 2.23 39.6 18.9 34.1 11.4
AuroraCap (DPOP [pal2024smaugfixingfailuremodes]) 1.55 1.88 1.78 1.79 1.96 2.30 41.1 19.3 34.8 11.5
AuroraCap (IPO [Azar2023AGT]) 1.44 1.83 1.74 1.76 1.95 2.21 39.8 19.1 34.3 11.3
AuroraCap (KTO [Ethayarajh2024KTOMA]) 1.53 1.88 1.73 1.78 1.97 2.27 40.1 19.0 34.3 11.4
AuroraCap (CPO [xu2024contrastive]) 1.42 1.81 1.75 1.72 1.94 2.25 40.8 18.9 34.1 11.2
AuroraCap (SimPO [meng2024simposimplepreferenceoptimization]) 1.52 1.83 1.76 1.75 1.96 2.26 40.2 19.1 34.5 11.3

### v2

@@ 1 Introduction
Fine-grained video captioning aims to generate detailed and coherent textual descriptions that precisely capture video contents. This task necessitates the recognition of salient actions and objects, while also modeling fine-grained visual features and temporal dynamics. Recent studies [undefn, undefaf, undefi, undefaq] have primarily employed Vision-Language Models (VLMs) for video captioning. These methods [undefaj, undefak, undefaaab] typically utilize pre-trained vision encoders and Large Language Models (LLMs), with a connector module linking them. By training on video-text pairs, these models aim to align visual and textual representations, thereby enhancing their ability to understand and describe video content [undefaao] effectively.
@@
Direct Preference Optimization (DPO) [undefaac] is a fine-tuning method that aligns models with stipulated preferences using high-quality preference pairs. Recent work has successfully adapted DPO and its variants to video understanding tasks, significantly improving model performance [undefan, undefal]. Thus, integrating DPO into fine-grained video captioning to enhance the model’s ability to capture temporal dynamics and detailed descriptions present a promising direction. However, two critical challenges currently degrade its performance in fine-grained video captioning: (1) the scarcity of high-quality video-text alignment pairs, which are essential for preference learning; (2) DPO typically suffers from the simultaneous decrease in both positive and negative reward values [undefaaa, undefm], leading to a potential objective optimization deviation from focusing on generation quality to merely discriminating between preferences [undefx], as shown in Figure 1 (middle).
@@
Recently, the proposed VDC benchmark [undefj] is well-curated, but limited in scale and partially reliant on manual annotations. Other video captioning datasets, such as MSRVTT [undefaay], VATEX [undefaaq], MSVC [undefn], etc., typically provide overly brief captions, falling short in fine-grained video captioning. Besides, these datasets lack preference pairs, and thus cannot be prepared for DPO. To construct preference pairs, many existing methods [undefc, undefam] rely on a stronger VLM to score multiple outputs from the same prompt. While straightforward, this approach is impractical: small teams face prohibitive API costs, while developers of powerful models often lack access to a stronger scoring model. Some studies attempt to circumvent this limitation by generating negative preferences, such as through atypical item substitution [undefaav] or temporal perturbation [undefan]. However, these methods primarily focus on negative samples and fail to produce higher-quality positive preferences.
@@
We present an automated pipeline for constructing high-quality preference pairs for fine-grained video captioning. Given the same input, we generate multiple alternative outputs using a VLM. These candidates are then scored leveraging intrinsic properties of the VLM itself, such as its self-consistency [undef] and the enhanced ability to capture details in short videos, with limited assistance from an LLM. The top and bottom scores are selected as positive and negative preferences, respectively, forming our constructed preference dataset. Compared to existing approaches, our method achieves an optimal balance between cost efficiency and high-quality preference pair construction.
@@
Our experiments across multiple models and datasets demonstrate that our data construction pipeline can generate high-quality preference datasets with considerable generality. In addition to video captioning benchmarks (e.g., VDC [undefj], VDD [undefak], VATEX [undefaaq], MSR-VTT [undefaay]), we conduct comparisons of SynPO with various DPO variants [undefv, undefaax, undefaaa, undefd, undefax] across multiple NLP tasks, which include preference evaluation tasks (e.g., MT-Bench [undefaaad], AlpacaEval2 [undefap]) and downstream applications (e.g., tasks from the Huggingface Open LLM Leaderboard [undefh, undefy]). As shown in Figure 1 (left), extensive results indicate that SynPO significantly outperforms DPO and its variants.
@@
Figure 1: Left: SynPO significantly outperforms other methods in different models on VDC benchmark [undefj]. Middle: Language capability degradation occurs during the latter training stages in DPO. Training collapses and is biased towards maximazing positive-negative reward gap. Right: SynPO mitigates degradtion successfully and resolves the issue of optimization objectives shifting from language capability to ranking differentiation. Its performance significantly outperforms that of DPO.
@@ 2 Related Work
Video Captioning. Early video captioning methods employed template-based approaches [undefab] or RNN-based encoder-decoder frameworks [undefaan], but were limited in modeling long-range dependencies. Transformers [undefaal] [undeft] [undefag] [undefs] [undefaab] and vision-language pre-training significantly advanced the field, with models like CLIP4Caption [undefaaj] and SwinBERT [undefar] improving video-text alignment. Recent VLMs [undefn, undefaq] adapt multimodal architectures such as BLIP-2 [undefaj] and LLaVA [undefau] to video, though many struggle with temporal dynamics. Newer approaches like VideoLLaMA [undefaaab] and ChatVideo [undefaao] better model sequential structure for enhanced comprehension.
Reinforcement Learning and Preference Learning. Reinforcement Learning from Human Feedback (RLHF) [undefo] aligns LLMs with human preferences [undefah, undefay, undefaah] through supervised fine-tuning [undefaaae], reward modeling [undefz], and policy optimization, improving instruction-following [undefaz] and safety [undeff]. To simplify RLHF’s complexity, offline methods like DPO [undefaac] bypass explicit reward modeling, inspiring variants such as IPO [undefe], ORPO [undefac] and others [undefaaz, undefaaw, undefaat, undefaas].
@@ 3.1 Enhanced Model Inference
Contrastive decoding was initially proposed by [undefai] to reduce object hallucinations and subsequently improved by [undefaaac], who introduced a more efficient variant that contrasts logits from sparse frame samples with those from full sequences. In this work, the improved contrastive decoding method is adopted to suppress overconfidence in noisy or irrelevant features, thereby improving factual consistency and detail accuracy in generated captions.
The self-retrospective strategy, proposed by [undefa], operates outside the decoding process and enhances comprehension through iterative refinement. Feeding the model’s own outputs back into its input enables a form of retrospective reasoning [undefaw], allowing predictions to be refined in light of prior generations. We adapt this method to video captioning by using the initial caption as contextual input for subsequent refinement steps, enabling richer and more coherent descriptions.
@@ 3.2 Dataset Construction Pipeline
We propose an automated pipeline for constructing high-quality preference pairs specifically designed for fine-grained video captioning. The overall framework is designed to address the limitations of existing datasets, such as limited scale, insufficient detail in captions, and the lack of human-like preference annotations [undefj, undefaaq, undefn]. Our method leverages both the intrinsic properties of VLMs and the reasoning capabilities of LLMs to generate diverse and reliable preference pairs without dependence on costly multimodal scorers.
The core strategy involves generating multiple candidate captions per video using a single VLM under the same prompt. These candidates are then scored using a novel three criterion evaluation framework that combines factuality, modality correctness, and self-consistency [undef]. Based on the aggregated scores across all three criteria, the captions with the highest and lowest total scores are selected as positive and negative preferences, respectively, forming our final dataset. In the following, we describe each scoring criterion in detail (full prompts provided in Appendix D).
Criterion 1: Factuality through Temporal Decomposition. Due to input length limits in most VLMs, processing long videos directly often causes detail loss and hallucinations [undefan]. To mitigate this, we divide each video into short clips, process them independently with the VLM to generate clip-level captions, and concatenate these into a reference set. An LLM then assesses the consistency between the full-video caption and the reference set, focusing particularly on factual alignment. This approach enhances detail preservation and mitigates hallucinations. Scores range from 0 to 5.
@@
Criterion 3: Self-consistency through Multi-sample Analysis. Inspired by self-consistency methods in NLP [undefaad, undefav, undefaaf], we apply it to video captioning by assessing the stability of key entities, actions, and temporal dynamics across multiple generations. Specifically, the VLM generates $n$ diverse captions via high-temperature sampling. An LLM analyzes their similarity, rewarding consistent patterns and penalizing outliers through a majority voting mechanism. Given its narrower discriminative capacity compared to the other two criteria, this metric uses a 0 to 3 scoring range.
@@ 4.1 Preliminary
In contrast, DPO [undefaac] introduces a simplified framework for preference optimization that bypasses explicit reward modeling. It directly formulates the preference optimization problem as a classification task over preference pairs, eliminating the need for a separate reward model. Compared with RLHF, DPO simplifies the training pipeline while achieving competitive performance in multiple tasks such as dialogue generation. Its objective function is defined as:
@@ 4.2.1 Existing Limitations
Despite practical success of DPO, several studies have identified limitations within DPO. Specifically, [undefaaa] point out that the DPO loss depends solely on the difference between the log-probability ratios of positive and negative preferences. Such a manner suggests that the final loss can decrease even when both are reduced, as long as the negative response decreases more rapidly. [undefax] observe preference optimization algorithms tend to decrease downstream task performance. Moreover, [undefaau] show that DPO frequently leads to imbalanced updates between positive and negative preferences. As illustrated in Figure 3, this imbalance manifests as a concurrent decline in both positive and negative rewards during training, particularly under higher learning rates.
@@ 4.2.2 Theoretical Insights
We argue that this substitution in the DPO derivation induces a fundamental deviation in the model’s optimization objective. In DPO, minimizing the loss is equated with improving the model’s ability to rank positive preferences above negative ones, rather than generating higher-quality outputs. Consequently, the model may behave more like a ranking model than a generative one. This deviation from the original goal of RLHF, namely, generating high-reward coherent text, can lead to suboptimal outcomes in terms of language capability. Furthermore, the logarithmic term derived from the KL-divergence constraint in Eq. (2) fails to serve its intended role in DPO. Due to the derivative properties of the logarithmic function, decreasing reward values require smaller gradient steps than increasing them [undefaau]. As a result, the optimizer is incentivized to reduce both positive and negative rewards simultaneously to minimize the overall loss.
@@ 4.2.3 Empirical Observations
To further investigate this behavior, we analyze the gradient dynamics of the DPO loss with respect to the policy parameters. According to the original DPO paper [undefaac], the gradient is derived as:
@@
However, empirical success has been reported in the original DPO paper and its variants (e.g., DPOP [undefaaa], IPO [undefd]).These results appear inconsistent with our theoretical findings on the limitations of DPO-style optimization. To understand this discrepancy, we analyze their experimental setups and identified two key factors contributing to the observed performance improvements:
(1) Low Learning Rates Mitigate Instability. Most DPO-style methods use significantly lower learning rates compared to standard Supervised Fine-Tuning (SFT). As shown in the original DPO paper [undefaac], a learning rate of 1e-6 is used—much lower than the typical SFT setting of 2e-5 (other variant configurations provided in Tabel 8). Our gradient-based analysis reveals that under such settings, the magnitude of parameter updates is less than one-tenth of that in standard SFT. This implicitly constrains the model’s deviation from its initial state, thereby alleviating the negative effects arising from the deviation of the optimization objective.
@@ 5.1 Constructing Long Video Caption Preference Pairs
Enhanced Inference Evaluation. The impact of contrastive decoding and the self-retrospective strategy on the quality of generated captions is evaluated. We compare four settings: baseline (no enhancement), contrastive decoding only, self-retrospective only, and the combination of both. Each setting is applied to generate captions on the dataset of VDD [undefak], and the outputs are evaluated by an LLM along six dimensions: accuracy, richness, completeness, fluency, dynamics and coherence (full prompts provided in Appendix D). As shown in Table 1, combining both strategies achieves the highest scores across all metrics, outperforming either method alone. Specifically: (1) Contrastive decoding significantly improves accuracy (+6.7%) and completeness (+2.8%), indicating its effectiveness in reducing hallucinations and ensuring factual consistency; (2) The self-retrospective strategy excels in enhancing richness (+5.5%) and dynamics (+7.6%), demonstrating its ability to inject more detailed content; (3) The combined approach retains the strengths of both methods, achieving balanced improvements in all aspects.
@@
To assess the contribution of each of our three proposed scoring criteria to the quality of preference pairs, we conduct ablation studies using AuroraCap [undefj] as the base model, generating 10 samples per input with the self-retrospective strategy. Results for different criterion combinations are reported in Table 2. The results show that all three criteria meaningfully contribute
@@
Downstream Preference Learning. To validate the effectiveness of the proposed pipeline, AuroraCap is fine-tuned using several methods, including SFT, DPO [undefaac], and our SynPO. The preference pairs used for fine-tuning were generated from a subset of Sharegpt4video [undefk] through our automated pipeline, with additonal details provided in Appendix C. Experiments are further conducted across diverse datasets and model configurations. The resulting models are evaluated on standard video captioning benchmarks (MSR-VTT [undefaay], VATEX [undefaaq]) using CIDEr (C) [undefaam] and METEOR (M) [undefg] metrics. In addition, we assess performance on VDC [undefj] and VDD [undefak] benchmarks, which employ LLM-based evaluation to better measure linguistic richness and factual consistency through their longer, richer reference captions(see Appendix C for details). As shown in Table 3, all fine-tuned models significantly surpassed their baselines, confirming that our data construction pipeline yields high-quality data.
Table 3: Experimental evaluation on various models (AuroraCap [undefj], LLaVA1.6-7B-video [undefat], InterVL2-8B [undefaap]), datasets (Sharegpt4video [undefk], Charades [undefaag], Pandas-70M [undefl]), and fine-tuning approaches (SFT, DPO, SynPO and other variants).
@@
AuroraCap (DPO [undefaac]) 1.39 1.89 1.70 1.73 1.94 2.23 39.6 18.9 34.1 11.4
AuroraCap (DPOP [undefaaa]) 1.55 1.88 1.78 1.79 1.96 2.30 41.1 19.3 34.8 11.5
AuroraCap (IPO [undefd]) 1.44 1.83 1.74 1.76 1.95 2.21 39.8 19.1 34.3 11.3
AuroraCap (KTO [undefv]) 1.53 1.88 1.73 1.78 1.97 2.27 40.1 19.0 34.3 11.4
AuroraCap (CPO [undefaax]) 1.42 1.81 1.75 1.72 1.94 2.25 40.8 18.9 34.1 11.2
AuroraCap (SimPO [undefax]) 1.52 1.83 1.76 1.75 1.96 2.26 40.2 19.1 34.5 11.3